How much should you invest now at 4 1/2% simple annual interest in order to earn $180 interest in 8 months?

Answer by stanbon(75887)

How much should you invest now at 4 1/2% simple annual interest in order to earn $180 interest in 8 months?
-------------------
I = P*r*t
180 = P*0.045*(8/12)
180 = P*0.045*(2/3)
----
0.045P = (3/2)180
----
0.045P = 270
----
P = $6000.00
--------------
Cheers,
Stan H.
-------------